OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 10 zero 6 Any Real Number No Solution

OpenStudy (tkhunny):

Use the distributive property to expand the expression on the left. Let's see what you get.

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

i think its D

OpenStudy (tkhunny):

Why do you think that? Show your work. You have intermediate results. Let us see them! 4(x + 1) = 4x + 10 Expand the left side 4x + 4 = 4x + 10 Subtract 4x 4 = 10 Well, that makes no sense! There is no solution.

OpenStudy (anonymous):

I have another question Solve: 7(x − 1) = 5(x + 2) + 1 9 18 2 4

OpenStudy (tkhunny):

Well, show your work and we'll see if you wander off. Expand all the parentheses.

OpenStudy (anonymous):

is it 9

OpenStudy (tkhunny):

I don't see any intermediate results. How do I know you didn't just pull that out of a hat?

OpenStudy (texaschic101):

oh...we have a magician

OpenStudy (anonymous):

its in an assessment Im taking and I want to see if Im right before I turn it in...thats why im asking for peoples opinions

OpenStudy (texaschic101):

distribute through the parenthesis and what do you get ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

let me check.....hold on

OpenStudy (anonymous):

cuz if u replace all x's with 9, they would be equal to eachother...am i correct?

OpenStudy (texaschic101):

7(x - 1) = 5(x + 2) + 1 7x - 7 = 5x +10 + 1 7x - 5x = 11 + 7 2x = 18 x = 9 that is correct

OpenStudy (anonymous):

thx guys

OpenStudy (tkhunny):

You didn't answer my objection. If the answer drops out of the sky, sure you can demonstrate that it works. This method is very unreliable. You must demonstrate your ability to MAKE the answers appear.

OpenStudy (texaschic101):

it takes too long to just plug in all of your answer choices to see which one is correct. You need to learn the distribution and solve the equation. It is much easier that way.

OpenStudy (anonymous):

again, I already figured out the answer, I just wanted to know if I was correct

OpenStudy (tkhunny):

I understand that explanation. I have heard it many times. It remains unhelpful. Please read the Code of Conduct and get a better feel how we work around here. We want to see what you are doing and we want to help you do it better. This following example conversation is neither helpful to you nor in compliance with the Code of Conduct. Q: I got 6. Is that right? A: Yes. Q: Thanks. A: YW That's not what we want to do, here.

OpenStudy (texaschic101):

He was just checking to see if his answer was correct. If it would not have been correct, I would have just said " no, try again ".

OpenStudy (texaschic101):

or asked to see how he arrived at that answer.

OpenStudy (anonymous):

I already know how to work the problems but thank you...I have to log off now cuz my mom needs the computer...

OpenStudy (texaschic101):

have a nice evening turtleluver :)

OpenStudy (tkhunny):

There are lots of ways to proceed. The very most helpful, the way that is most beneficial to the student and doesn't waste any time from volunteers is this: Q: I have this problem. I think I have it right. This is what I did... I'm a little nervous about the second step. What do you think? A: Your second step is perfect. Pat yourself on the back, gain a little confidence and move on to the next problem. Alternatively: Q: I have this problem. I think I have it right. This is what I did... I'm a little nervous about the second step. What do you think? A: You were wise to suspect step 2. You should have combined those terms on the left side to get... Now, give it another go and see what you get on the next try. I may have limitless time to work algebra problems, but as a general rule, we want to encourage volunteers to keep volunteering. Making them blind answer checkers does not accomplish this worthwhile social goal.

OpenStudy (anonymous):

u too

OpenStudy (texaschic101):

I understand your point tkhunny. I will try to do better. Very good examples by the way :)